given h(x) = -x2, find h(-2/5)
Answer by AnlytcPhil(1116)

given h(x) = -x2, find h(-2/5)

Although you were originally taught that h(x) and h(-2/5)
mean "h times x" and "h times -2/5", it NO LONGER MEANS THAT!!!!
It now means to replace x by -2/5.

           h(x) = -x2

1. Take out the x on the left, leaving an open parentheses

          h(    ) = -x2

2. Take out all x's on the right, replacing them by a pair of
   empty open parentheses (    )

          h(    ) = -(    )2

3. Put -2/5 into the empty parentheses on both sides:

          h(-2/5) = -(-2/5)2

4. Next find (-2/5)2 by multiplying -2/5 times -2/5, getting +4/25, so
   replace (-2/5)2 by (4/25)

          h(-2/5) = -(4/25)

          h(-2/5) = -4/25

That's the answer.

Edwin
AnlytcPhil@aol.com

h(-2/5) = -4/25